OpenStudy (anonymous):

A teacher set up groups of three students to work together. She asked them to solve the equation independently, then check their answers with each other. Here are the results: x divided by three minus two-fifths equals negative one-fifth Chet: x equals negative three-fifths Lauren: x equals one and four-fifths Bradley: x equals three-fifths Which student found with the correct answer? Chet Lauren Bradley None of them

OpenStudy (yanasidlinskiy):

x/3 - 2/5 = -1/5 --- add 2/5 to each side x/3 = -1/5 + 2/5 --- combine like terms x/3 = 1/5 --- multiply entire equation by 3 x = 1/5(3) x = 3/5 Which would make Bradley correct!

OpenStudy (anonymous):

omg thank youuuu :):):):):):):):):):):):):):):):):)

OpenStudy (yanasidlinskiy):

No problem!!:) Your welcome! Anytime!!!:)